A girl inside of a large inflatable ball (called a Zorbing) starts rolling down a grassy hill at
an acceleration of 3.6 m/s^2. The hill is 12 m in length. How fast is she going at the
bottom of the hill?
1
Expert's answer
2020-12-04T07:33:06-0500
"v^2=2as\\v^2=2(3.6)(12)\\v=9.3\frac{m}{s}"